### v3

@@ 5.3 Implementation
We provide a free and open source implementation of our proposed methodology in a package for the statistical programming environment R (R Core Team,, 2024). The package is called robcat (for “ROBust CATegorical data analysis”) and is available at https://github.com/mwelz/robcat. To maximize speed and performance, the package is predominantly developed in C++ and integrated to R via Rcpp (Eddelbuettel,, 2013). All numerical results in this paper were obtained with this package.
The estimator’s minimization problem in (10) can be solved with standard algorithms for numerical optimization. Yet, for the robust estimator we recommend to use constrained optimization techniques for imposing the boundary constraint on the correlation coefficient and the monotonicity constraints on the thresholds; see (5). In package robcat, the simplex algorithm of Nelder & Mead, (1965) for constrained optimization is used for this purpose.
An important user choice is that of the tuning constant $c$ in discrepancy function (11). The closer $c$ is to 0, the more robust the estimator will be to possible misspecification of the polychoric model (see, e.g., Figure 3). On the other hand, in the presence of model misspecification, the more robust the estimator is made, the larger its estimation variance becomes. Moreover, if the model is correctly specified, then Welz, (2024) shows that the most robust choice, $c=0$, is associated with two drawbacks, namely asymptotic nonnormality as well as certain finite sample issues. To avoid these drawbacks, we suggest to choose a value slightly larger than 0. In simulation experiments, we found that the choice $c=0.6$ yields a good compromise between robustness and efficiency in case of misspecification while not suffering from finite sample issues if the model is correctly specified. We therefore chose to use this value for all applications in this paper. However, we acknowledge that a detailed study, preferably founded in statistical theory, is necessary to provide guidelines on the choice of $c$. We will explore this in future work.
@@ 6 Simulation studies on partial misspecification
In this section, we employ two simulation studies to demonstrate the robustness gain of our proposed estimator under partial misspecification of the polychoric model. The first simulation design is a simplified setting with respect to the partial misspecification, chosen specifically to illustrate the effects of a particular misspecification type with high leverage affecting only a small number of contingency table cells. The second design is more involved and considers estimation of a polychoric correlation matrix with a misspecification type that scatters in many directions so that nonnormal data points can occur in every contingency table cell.
@@ 6.1 Individual polychoric correlation coefficient
and let the true polychoric correlation coefficient in latent normality model (2) be $\rho_{*}=0.5$. To simulate partial misspecification of the polychoric model, we let a fraction $\varepsilon$ of the data be generated by a misspecifying bivariate normal distribution with mean $(2,-2)^{\top}$, variances $(0.2,0.2)^{\top}$, and zero covariance (and therefore zero correlation). This misspecifying distribution will inflate the empirical frequency of contingency table cells $(x,y)\in\{(5,1),(4,3),(5,2)\}$, in the sense that they have a higher realization probability than under the true polychoric model. In fact, the data plotted in Figure 1 were generated by this process for misspecification fraction $\varepsilon=0.15$, and one can see in this figure that particularly cell $(x,y)=(5,1)$ is sampled quite frequently although it only has a near-zero probability at the true polychoric model. The data points causing these three cells to be inflated are instances of negative leverage points. Here, such leverage points drag correlational estimates away from a positive value towards zero or, if there are sufficiently many of them, even a negative value. One may think of such points as the responses of careless or inattentive participants whose responses are not based on item content.
@@
For misspecification fraction $\varepsilon\in\{0,0.01,0.05,0.1,0.15,0.2\}$, we sample $N=1,000$ ordinal responses from this data generating process and estimate the true parameter $\bm{\theta}_{*}$ with the MLE (Olsson,, 1979), our proposed estimator with tuning constant set to $c=0.6$, and, for comparison, the Pearson sample correlation calculated on the integer-valued responses. We repeat this procedure for 1,000 simulated datasets. As performance measures, we calculate the average bias, standard deviation across repetitions, coverage, and length of confidence intervals at significance level $\alpha=0.05$. The coverage is defined as proportion of $(1-\alpha)$-th confidence intervals $[\widehat{\rho}_{N}\mp q_{1-\alpha/2}\cdot\text{SE}(\widehat{\rho}_{N})]$ that contain the true $\rho_{*}$, where $q_{1-\alpha/2}$ is the $(1-\alpha/2)$-th quantile of the standard normal distribution and $\text{SE}(\widehat{\rho}_{N})$ is the standard error of $\widehat{\rho}_{N}$, which is constructed using the limit theory developed in Theorem 1 in Appendix A. The length of a confidence interval is given by $2\cdot q_{1-\alpha/2}\cdot\text{SE}(\widehat{\rho}_{N})$.
@@
Figure 4: Boxplot visualization of the bias of three estimators of the polychoric correlation coefficient, $\widehat{\rho}_{N}-\rho_{*}$, for various degrees of misspecification of the polychoric model across 1,000 replications. The estimators are the proposed robust estimator with $c=0.6$ (left), the MLE of Olsson, (1979) (center), and the Pearson sample correlation coefficient (right).
Misspecification Estimator $\widehat{\rho}_{N}$ Bias StDev Coverage CI length
$\varepsilon=0$ Polycor robust 0.505 0.005 0.027 0.942 0.104
Polycor ML 0.501 0.001 0.027 0.949 0.102
Sample cor 0.457 $-0.043$ 0.025 0.686 0.110
$\varepsilon=0.01$ Polycor robust 0.497 $-0.003$ 0.029 0.944 0.110
Polycor ML 0.447 $-0.053$ 0.027 0.528 0.109
Sample cor 0.408 $-0.092$ 0.025 0.073 0.113
$\varepsilon=0.05$ Polycor robust 0.479 $-0.021$ 0.033 0.912 0.130
Polycor ML 0.272 $-0.228$ 0.029 0.000 0.126
Sample cor 0.242 $-0.258$ 0.026 0.000 0.120
$\varepsilon=0.1$ Polycor robust 0.466 $-0.034$ 0.039 0.906 0.153
Polycor ML 0.096 $-0.404$ 0.029 0.000 0.134
Sample cor 0.083 $-0.417$ 0.026 0.000 0.124
$\varepsilon=0.15$ Polycor robust 0.456 $-0.044$ 0.044 0.922 0.176
Polycor ML $-0.050$ $-0.550$ 0.029 0.000 0.136
Sample cor $-0.038$ $-0.538$ 0.02y 0.000 0.124
$\varepsilon=0.2$ Polycor robust 0.440 $-0.060$ 0.052 0.943 0.220
Polycor ML $-0.173$ $-0.673$ 0.028 0.000 0.133
Sample cor $-0.134$ $-0.634$ 0.026 0.000 0.123
@@
Table 1: Performance measures of of the three estimators, Pearson sample correlation, the MLE, and the proposed robust estimator with $c=0.6$, for various misspecification fractions across 1,000 simulated datasets. The true polychoric correlation coefficient is $\rho_{*}=0.5$. The performance measures are the average point estimate of the polychoric correlation coefficient, $\widehat{\rho}_{N}$, average bias ($\widehat{\rho}_{N}-\rho_{*}$), the standard deviation of the $\widehat{\rho}_{N}$ (“StDev”), the estimator’s coverage with respect to the true $\rho_{*}$ at significance level $\alpha=0.05$, as well as the length of the estimator’s confidence interval, again at level $\alpha=0.05$.
@@
Figure 4 visualizes the bias of each estimator with respect to the true polychoric correlation $\rho_{*}$ across the  1,000 simulated datasets. An analogous plot for the whole parameter $\bm{\theta}_{*}$ can be found in Appendix C; the results are similar to those of $\rho_{*}$. For all considered misspecification fractions, the estimates of the MLE and sample correlation are similar, which is expected because these two estimators are known to yield similar results when there are five or more rating options and the discretization thresholds are symmetric (cf., Rhemtulla et al.,, 2012). In the absence of misspecification, both MLE and the robust estimator yield accurate estimates. Both estimates are nearly equivalent to one another in the sense that their point estimates, standard deviation, and coverage at significance level $\alpha=0.05$ are very similar (Table 1). However, when misspecification is introduced, MLE, sample correlation, and robust estimator yield noticeably different results. Already at the small misspecification fraction $\varepsilon=0.01$ (corresponding to only 10 observations), MLE and sample correlation are noticeably biased, resulting in poor coverage of only about 0.53 and 0.07, respectively. Increasing the misspecification fraction to the still relatively small value of $\varepsilon=0.05$, MLE and sample correlation start to be substantially biased with average biases of $-0.23$ and $-0.26$, respectively, leading to zero coverage. The biases of these two methods further deteriorates as the misspecification fraction is gradually increased. At $\varepsilon=0.2$, MLE and sample correlation produce estimates that are not only severely biased (biases of $-0.67$ and $-0.63$), but also sign-flipped: While the true correlation is strongly positive (0.5), both estimates are considerably negative ($-0.17$ and $-0.13$). In stark contrast, the proposed robust estimator remains accurate throughout all considered misspecification fractions: At the small $\varepsilon=0.01$, the robust estimator is nearly unaffected, while at the larger $\varepsilon=0.2$, it only exhibits a minor bias of about $-0.06$. In all considered scenarios, its coverage remained above 0.9. It is worth noting that in the presence of misspecification, the confidence intervals of the robust estimator are wider than those of the MLE (see Table 1). This is expected because of a well-known trade-off between robustness and efficiency: An estimator that is designed to reduce bias, like a robust estimator, will inevitably have a larger estimation variance (e.g., Huber & Ronchetti,, 2009). These wider confidence intervals furthermore explain why the robust estimator improves its coverage in Table 1 when the misspecification fraction is increased from 0.1 to 0.15 and 0.2.
@@ 6.2 Polychoric correlation matrix
Let there be $q$ observed ordinal random variables and assume that a latent variable underlies each ordinal variable. In accordance with the multivariate polychoric model (e.g., Muthén,, 1984), the latent variables are assumed to jointly follow a $q$-dimensional normal distribution with mean zero and a covariance matrix with unit diagonal elements so that the covariance matrix is a correlation matrix. Each individual latent variable is discretized to its corresponding observed ordinal variable akin to discretization process (1).
@@
Table 2: Correlation matrix of $q=5$ latent variables as in Foldnes & Grønneberg, (2020). In line with the multivariate polychoric correlation model (e.g., Muthén,, 1984), the latent variables are jointly normally distributed with mean zero and this correlation matrix as covariance matrix.
Following the five-dimensional simulation design in Foldnes & Grønneberg, (2020), there are $q=5$ ordinal variables with polychoric correlation matrix as in Table 5 such that the pairwise correlations vary from a low 0.2 to a moderate 0.56.444The correlation matrix in Table 2 has the additional interpretation of being the covariance matrix of a factor model for a single factor with loadings vector $(0.8,0.7,0.6,0.5,0.4)^{\top}$. For all latent variables, the discretization thresholds are set to, in ascending order, $\Phi_{1}^{-1}(0.1)=-1.28,\ \Phi_{1}^{-1}(0.3)=-0.52,\ \Phi_{1}^{-1}(0.7)=0.56$, and $\Phi_{1}^{-1}(0.9)=1.28$, such that each ordinal variable can take five possible values. A visualization of the implied distribution of each ordinal variable can be found in Figure 5 in Foldnes & Grønneberg, (2020).
As misspecifying distribution, we choose a $q$-dimensional Gumbel distribution comprising of mutually independent Gumbel marginal distributions, each with location and scale parameters equal to 0 and 3, respectively. To obtain ordinal observations, the unobserved realizations from this distribution are discretized via the same threshold values as realizations from the model (normal) distribution. As such, the uninformative ordinal observations generated by this misspecifying distribution emulate the erratic behavior of a careless respondent: Unlike in the previous simulation design (Section 6.1), the uninformative responses are not concentrated around a few response options, but may occur in every response option.
For misspecification fraction $\varepsilon\in\{0,0.01,0.05,0.1,0.15,0.2\}$, we sample $N=1,000$ ordinal five-dimensional responses from this data generating process and use them to estimate the polychoric correlation matrix in Table 2 via ML and our robust estimator (again with tuning constant $c=0.6$). We repeat this procedure 1,000 times.
@@
Figure 5: Absolute average bias of the ML estimator (top) and our robust estimator (bottom) for each unique pairwise polychoric correlation coefficient in the true correlation matrix (Table 2), expressed as a function of the misspecification fraction $\varepsilon$ ($x$-axis). Averages are taken over 1,000 repetitions of the data generating process described in Section 6.2.
@@
Figure 5 visualizes the absolute average bias (calculated over the 1,000 repetitions) of the ML and robust estimator for each pairwise polychoric correlation coefficient in Table 2. As expected, when the model is correctly specified ($\varepsilon=0$), then both estimators coincide with accurate estimates. However, in the presence of partial misspecification ($\varepsilon>0$), the two estimators deviate. The ML estimator exhibits a notable bias for all correlation coefficient, which increases with increasing misspecification fraction. The magnitude of the bias tends to be larger for pairs with larger true correlation, such as $0.56$ for pair $(1,2)$, than for pairs with weaker true correlation, such as $0.20$ for pair $(5,4)$. Conversely, the robust estimator remains nearly unaffected throughout all considered misspecification fractions for each correlation coefficient, reflecting an excellent performance with respect to robustness to uninformative responses.

### v4

@@ 5.3 Implementation
We provide a free and open source implementation of our proposed methodology in a package for the statistical programming environment R (R Core Team,, 2024). The package is called robcat (for “ROBust CATegorical data analysis”; Welz et al.,, 2025), and it is available from CRAN (the Comprehensive R Archive Network) at https://CRAN.R-project.org/package=robcat. To maximize speed and performance, the package is predominantly developed in C++ and integrated to R via Rcpp (Eddelbuettel,, 2013). All numerical results in this paper were obtained with this package.
The estimator’s minimization problem in (10) can be solved with standard algorithms for numerical optimization. In our experience, using an unconstrained version of the quasi-Newton algorithm L-BFGS-B of Byrd et al., (1995) works fine. However, additional stability might be gained from imposing the boundary constraint on the correlation coefficient and the monotonicity constraints on the thresholds, see (5), for which the the simplex algorithm of Nelder & Mead, (1965) for constrained optimization can be used. In our implementation in package robcat, the default behavior is to first try unconstrained optimization via L-BFGS-B. If numerical instability is encountered or a monotonicity constraint is violated, the constrained optimization algorithm of Nelder & Mead, (1965) is used instead. While this is the default behavior, the package allows users to freely specify any supported optimization routine.
An important user choice is that of the tuning constant $c$ in discrepancy function (11). The closer $c$ is to 0, the more robust the estimator will be to possible misspecification of the polychoric model (see, e.g., Figure 3). On the other hand, in the presence of model misspecification, the more robust the estimator is made, the larger its estimation variance becomes. Moreover, if the model is correctly specified, then Welz, (2024) shows that the most robust choice, $c=0$, is associated with two drawbacks, namely asymptotic nonnormality as well as certain finite sample issues. We therefore suggest choosing a value slightly larger than 0. In simulation experiments (see Appendix C), we find that the estimator is relatively insensitive to the specific choice of $c>0$, as long as it is reasonably small (for robustness) yet sufficiently far away from 0 (to avoid the aforementioned issues). The choice $c=0.6$ thereby yields a good compromise so that we use this value for all applications in this paper. However, we acknowledge that a detailed study, preferably founded in statistical theory, is necessary to provide guidelines on the choice of $c$. We will explore this in future work.
@@ 6 Simulation studies on partial misspecification
In this section, we employ two simulation studies to demonstrate the robustness gain of our proposed estimator under partial misspecification of the polychoric model. The first simulation design (Section 6.1) is a simplified setting with respect to the partial misspecification, chosen specifically to illustrate the effects of a particular type of contamination with high leverage affecting only a small number of contingency table cells. The second design (Section 6.2) is more involved and considers estimation of a polychoric correlation matrix with a contamination type that scatters in many directions so that nonnormal data points can occur in every contingency table cell. Section 6.3 summarizes findings from additional simulations in the appendix. For all simulation designs, we perform 5,0005,0005 , 000 repetitions.
@@ 6.1 Individual polychoric correlation coefficient
[…]
For contamination fraction $\varepsilon\in\{0,0.01,0.05,0.1,0.15,0.2,0.3,0.4,0.49\}$, we sample $N=1,000$ ordinal responses from this data generating process. We estimate the true parameter $\bm{\theta}_{*}$ with our proposed estimator with tuning constant set to $c=0.6$, the MLE (Olsson,, 1979), and, for comparison, the Pearson sample correlation calculated on the integer-valued responses.
@@
Let $\widehat{\rho}_{N}$ be the estimate on a simulated dataset and $\widehat{\textnormal{SE}}\left(\widehat{\rho}_{N}\right)$ the estimated standard error of $\widehat{\rho}_{N}$, which is constructed using the limit theory developed in Theorem 1 in Appendix A. As performance measures, we calculate the average bias of the correlation estimates, the average bias of the standard error estimates (using the standard deviation of the correlation estimates across repetitions as an approximation of the true standard error), as well as coverage and average length of confidence intervals at significance level $\alpha=0.05$. The coverage is defined as the proportion (across repetitions) of confidence intervals $\left[\widehat{\rho}_{N}\mp q_{1-\alpha/2}\cdot\widehat{\textnormal{SE}}\left(\widehat{\rho}_{N}\right)\right]$ that contain the true $\rho_{*}$, where $q_{1-\alpha/2}$ is the $(1-\alpha/2)$ quantile of the standard normal distribution. The length of a confidence interval is given by $2\cdot q_{1-\alpha/2}\cdot\widehat{\textnormal{SE}}\left(\widehat{\rho}_{N}\right)$.
@@
Figure 4: Boxplot visualization of the bias of three estimators of the polychoric correlation coefficient, $\widehat{\rho}_{N}-\rho_{*}$, for various contamination fractions in the misspecified polychoric model across 5,000 repetitions. The estimators are the robust estimator with $c=0.6$ (left), the MLE (center), and the Pearson sample correlation (right). Diamonds represent the respective average bias. The dashed line denotes value 0 and the dotted line $-\rho_{*}=-0.5$, the latter of which indicating a sign flip in the correlation estimate.
Point estimate Standard error Confidence interval
Contamination Estimator $\widehat{\rho}_{N}$ Bias SE $\widehat{\text{SE}}$ Bias Coverage Length
$\varepsilon=0$ Robust 0.504 0.004 0.027 0.027 0.001-0.001- 0.001 0.931 0.104
ML 0.500 0.000 0.027 0.026 0.001-0.001- 0.001 0.939 0.102
Sample cor 0.456 0.044-0.044- 0.044 0.025 0.025 0.003 0.690 0.110
$\varepsilon=0.01$ Robust 0.501 0.001 0.028 0.027 0.001-0.001- 0.001 0.938 0.107
ML 0.441 0.059-0.059- 0.059 0.027 0.028 0.001 0.446 0.110
Sample cor 0.402 0.098-0.098- 0.098 0.025 0.029 0.004 0.043 0.114
$\varepsilon=0.05$ Robust 0.495 0.005-0.005- 0.005 0.029 0.029 0.001-0.001- 0.001 0.939 0.112
ML 0.252 0.248-0.248- 0.248 0.028 0.033 0.005 0.000 0.128
Sample cor 0.221 0.279-0.279- 0.279 0.025 0.031 0.006 0.000 0.121
$\varepsilon=0.1$ Robust 0.488 0.012-0.012- 0.012 0.031 0.031 0.000 0.933 0.120
ML 0.062 0.438-0.438- 0.438 0.028 0.035 0.006 0.000 0.135
Sample cor 0.052 0.448-0.448- 0.448 0.025 0.032 0.007 0.000 0.124
$\varepsilon=0.15$ Robust 0.480 0.020-0.020- 0.020 0.033 0.033 0.000 0.907 0.130
ML 0.097-0.097- 0.097 0.597-0.597- 0.597 0.027 0.035 0.007 0.000 0.135
Sample cor 0.076-0.076- 0.076 0.576-0.576- 0.576 0.025 0.032 0.007 0.000 0.124
$\varepsilon=0.2$ Robust 0.472 0.028-0.028- 0.028 0.036 0.036 0.001 0.882 0.142
ML 0.229-0.229- 0.229 0.729-0.729- 0.729 0.026 0.033 0.008 0.000 0.131
Sample cor 0.176-0.176- 0.176 0.676-0.676- 0.676 0.025 0.031 0.007 0.000 0.122
$\varepsilon=0.3$ Robust 0.450 0.050-0.050- 0.050 0.043 0.045 0.002 0.789 0.176
ML 0.431-0.431- 0.431 0.931-0.931- 0.931 0.022 0.030 0.008 0.000 0.118
Sample cor 0.322-0.322- 0.322 0.822-0.822- 0.822 0.024 0.030 0.007 0.000 0.117
$\varepsilon=0.4$ Robust 0.413 0.087-0.087- 0.087 0.053 0.067 0.014 0.595 0.261
ML 0.571-0.571- 0.571 1.071-1.071- 1.071 0.019 0.026 0.008 0.000 0.103
Sample cor 0.424-0.424- 0.424 0.924-0.924- 0.924 0.024 0.029 0.005 0.000 0.112
$\varepsilon=0.49$ Robust 0.357 0.143-0.143- 0.143 0.062 0.071 0.009 0.247 0.278
ML 0.662-0.662- 0.662 1.162-1.162- 1.162 0.016 0.024 0.008 0.000 0.095
Sample cor 0.492-0.492- 0.492 0.992-0.992- 0.992 0.023 0.028 0.004 0.000 0.108
@@
Table 1: Results for the robust estimator with $c=0.6$, the MLE, and the Pearson sample correlation, for various contamination fractions across 5,000 simulated datasets. The true polychoric correlation coefficient is $\rho_{*}=0.5$. We compute the average of point estimates $\widehat{\rho}_{N}$ of the polychoric correlation coefficient, the average bias ($\widehat{\rho}_{N}-\rho_{*}$), the standard deviation of the $\widehat{\rho}_{N}$ (SE; an approximation of the true standard error), the average standard error estimate $\widehat{\text{SE}}$, the corresponding average bias ($-\text{SE}$ SE), confidence interval coverage with respect to the true $\rho_{*}$ at nominal level 95%, and the average length of the respective confidence intervals.
@@
Figure 4 visualizes the bias of each estimator with respect to the true polychoric correlation $\rho_{*}$ across the 5,000 simulated datasets. An analogous plot for the whole parameter $\bm{\theta}_{*}$ can be found in Appendix D.1; the results are similar to those of $\rho_{*}$. Additional performance measures are shown in Table 1. For all considered contamination fractions, the estimates of the MLE and sample correlation are somewhat similar, which is expected because these two estimators are known to yield similar results when there are five or more rating options and the discretization thresholds are symmetric (cf., Rhemtulla et al.,, 2012). In the absence of contamination, the MLE and the robust estimator yield accurate estimates. Both estimators are nearly equivalent to one another in the sense that their point estimates, standard deviation, and coverage at significance level $\alpha=0.05$ are very similar. However, when contamination is introduced, MLE, sample correlation, and the robust estimator yield noticeably different results. Already at the small contamination fraction $\varepsilon=0.01$ (corresponding to only 10 observations), MLE and sample correlation are noticeably biased, resulting in poor coverage of only about 0.45 and 0.04, respectively. Increasing the contamination fraction to the still relatively small value of $\varepsilon=0.05$, MLE and sample correlation start to be substantially biased with average biases of 0.25-0.25- 0.25 and 0.28-0.28- 0.28, respectively, leading to zero coverage. The biases of these two methods further deteriorate as the contamination fraction is gradually increased. At $\varepsilon\geq 0.15$, MLE and sample correlation produce estimates that are not only severely biased but also sign-flipped: while the true correlation is positive (0.5), both estimates are negative. In stark contrast, the proposed robust estimator remains accurate throughout nearly all considered contamination fractions. At the small $\varepsilon=0.01$, the robust estimator is nearly unaffected, while at $\varepsilon=0.15$, it only exhibits a minor bias of about 0.02-0.02- 0.02. Even at extreme contamination $\varepsilon=0.4$, its bias amounts to less than 0.1. In addition, coverage of the robust method remains above or close to 0.9 for contamination fractions $\varepsilon\leq 0.2$.
It is worth noting that the confidence intervals of the robust estimator grow wider with increasing contamination fraction $\varepsilon$. We also observe that the standard deviation of the robust estimates over the repetitions grow similarly. This indicates that the derived asymptotic distribution used to estimate standard errors matches well with the simulated distribution of the estimator across the repetitions. Indeed, the bias of standard error estimation remains at near-zero for the robust estimator, except for extremely large contamination fractions ($\varepsilon\geq 0.4$). We investigate this in more detail in Appendix D.1.
@@ 6.2 Polychoric correlation matrix
Let there be $r$ observed ordinal random variables and assume that a latent variable underlies each ordinal variable. In accordance with the multivariate polychoric model (e.g., Muthén,, 1984), the latent variables are assumed to jointly follow an $r$-dimensional normal distribution with mean zero and a covariance matrix with unit diagonal elements so that the covariance matrix is a correlation matrix. Each individual latent variable is discretized to its corresponding observed ordinal variable akin to discretization process (1).
Following the five-dimensional simulation design in Foldnes & Grønneberg, (2020), there are $r=5$ ordinal variables with polychoric correlation matrix as in Table 2 such that the pairwise correlations vary from a low 0.2 to a moderate 0.56.121212The correlation matrix in Table 2 has the additional interpretation of being the covariance matrix of a factor model for a single factor with loadings vector $(0.8,0.7,0.6,0.5,0.4)^{\top}$. For all latent variables, the discretization thresholds are set to, in ascending order, $\Phi_{1}^{-1}(0.1)=-1.28,\ \Phi_{1}^{-1}(0.3)=-0.52,\ \Phi_{1}^{-1}(0.7)=0.56$, and $\Phi_{1}^{-1}(0.9)=1.28$, such that each ordinal variable can take five possible values. A visualization of the implied distribution of each ordinal variable can be found in Figure 5 in Foldnes & Grønneberg, (2020).
As contamination distribution, we choose an $r$-dimensional Gumbel distribution comprising of mutually independent Gumbel marginal distributions, each with location and scale parameters equal to 0 and 3, respectively. To obtain ordinal observations, the unobserved realizations from this distribution are discretized via the same threshold values as realizations from the model (normal) distribution. As such, the uninformative ordinal observations generated by this contaminated distribution emulate the erratic behavior of a careless respondent. Unlike in the previous simulation design (Section 6.1), the uninformative responses are not concentrated around a few response options, but may occur in every response option.
For contamination fraction $\varepsilon\in\{0,0.01,0.05,0.1,0.15,0.2,0.3,0.4,0.49\}$, we sample $N=1,000$ ordinal five-dimensional responses from this data generating process and use them to estimate the polychoric correlation matrix in Table 2 via our robust estimator (again with tuning constant $c=0.6$) as well as the MLE.
@@
Table 2: Correlation matrix of $r=5$ latent variables as in Foldnes & Grønneberg, (2020). In line with the multivariate polychoric correlation model (e.g., Muthén,, 1984), the latent variables are jointly normally distributed with mean zero and this correlation matrix as covariance matrix.
@@
Figure 5: Absolute average bias (top) and confidence interval coverage (bottom) at nominal level 95% (dashed horizontal lines) of the robust estimator with $c=0.6$ (left) and the MLE (right) for each unique pairwise polychoric correlation coefficient in the true correlation matrix (Table 2), expressed as a function of the contamination fraction $\varepsilon$ ($x$-axis). Results are aggregated over 5,000 repetitions.
@@
Figure 5 visualizes the absolute average bias as well as confidence interval coverage at 95% nominal level (calculated over the 5,000 repetitions) of the robust estimator and the MLE for each pairwise polychoric correlation coefficient. As expected, when the model is correctly specified ($\varepsilon=0$), both estimators coincide with accurate estimates. However, in the presence of contamination ($\varepsilon>0$), the two estimators deviate. The MLE exhibits a notable bias for all correlation coefficients, which increases gradually with increasing contamination fraction. The magnitude of the bias tends to be larger for pairs with larger true correlation, such as 0.560.560.56 for pair (1,2)(1,2)( 1 , 2 ), than for pairs with weaker true correlation, such as 0.200.200.20 for pair (5,4)(5,4)( 5 , 4 ). In addition, coverage of the MLE drops quickly for many pairs and gradually for the remaining ones. Conversely, the robust estimator remains nearly unaffected for a broad range of contamination fractions for each correlation coefficient ($\varepsilon\leq 0.2$ or $\varepsilon\leq 0.3$ depending on the variable pair), with bias only somewhat increasing afterwards. Furthermore, its coverage remains close to 0.9 or higher even at the high contamination fraction of $\varepsilon=0.3$. This reflects excellent performance of the proposed estimator with respect to robustness to uninformative responses. Appendix D.2 contains additional evaluations. In essence, the robust estimator yields accurate standard errors, but its confidence intervals tend to be wider than those of the MLE in the presence of contamination.
6.3 Discussion and additional simulation experiments